07/09/08 – TTC stopped by the site today and review the project progress with Process. There were several issues with the installation of the equipment that we resolved. Size, location, pipe runs, housekeeping pad, inlets.

12/04/07 – We received a schedule of values today that TTC marked up and sent back to Process. Insulation submittals were received today in the mail. LKV is logging them in for this project and will give them to TTC to review.

11/29/07 – Joe Doxsey and Terry Cannon discussed project schedule and submittals today by phone. Process is going to send their background checks to Willie at the County and are currently working on the submittals. They have not worked on submittals or project schedule yet according to Joe but will get them done now and into the system as discussed with TEI today.

11/16/07 – We held the pre construction meeting with Process and discussed changed scope and timetable for the project. Bert provided an email with the PO attached yesterday. TTC asked for the submittals on the vacuum pumps for the job and saw Rick Huff today and asked him if they were ready. He had no idea about them being asked for so Process must not have told him yet. Laurie finished typing the meeting notes and TTC has to review them and send them out. Process is going to provide a project schedule and SOV for our approval and ThermalTech has printed out and is sending 9 sets of stamped permit drawings (6 for city, 3 for IBI) for Process use on this job. TEI also emailed Process a copy of the spec and the dwgs for their personal use and printing.

08/02/07 – TTC sent out Addendum 02 for the project today PDF. Addendum 01 extended the bid opening to August 7 by Purchasing.

07/31/07 – Bernie and Roger replied with Addendum 01 items for this project. A&M stated that their cost would be about $4500 for this work even though we are adding about 8 points and picking up alarms for the system. Bernie added that the 3″ line sizes were all he needed from the dwgs.

07/30/07 – Purchasing issued Addendum 01 for this project extending the Bid Opening and Due Date to August 7, 2007. TTC will issue Addendum 2 with pipe sizes, control notes and spec revisions.

07/16/07 – TTC sent the same data to Queen City to hold for bidders this morning. TTC has to contact Gina about the advertised price for this project since he changed it after talking to Tony Matre.

07/13/07 – TTC sent the project out for bidding with a bid opening date of July 31, 2007. Rod, Jim and Michael helped complete the project and send it to Gina by 4pm.

06/26/07 – TTC and Rod are going to finish the mechanical design on this project. Michael Huntwork is going to finish the electrical work on the project and TTC has to get project number from Purchasing for this project.

05/30/07 – TTC selected pumps and is finalizing the dwgs and specs for bidding. TTC asked to add the dampers to the back of the combustion air intake while A&M was doing the control work on this job.

05/14/07 – TTC and NR worked on CAD layout and background drawings. TTC will try to have something to review by this Friday. TTC talked to Tony today about using 2 full size vacuum pumps instead of 3 separate pumps. If this works, Tony did not seem to have a problem with the new design.

05/03/07 – TTC visited the site for field work and equipment locations. Meeting notes were distributed to Owner.

04/13/07 – We had second Design Mtg. 3 vendors were selected. We filled out the BOD and DID for the project.

03/15/07 – Tony said the SOSS is approved but TE has not seen it yet.